# Contrastive Difference Predictive Coding

### v2

@@ Abstract
Predicting and reasoning about the future lie at the heart of many time-series questions. For example, goal-conditioned reinforcement learning can be viewed as learning representations to predict which states are likely to be visited in the future. While prior methods have used contrastive predictive coding to model time series data, learning representations that encode long-term dependencies usually requires large amounts of data. In this paper, we introduce a temporal difference version of contrastive predictive coding that stitches together pieces of different time series data to decrease the amount of data required to learn predictions of future events. We apply this representation learning method to derive an off-policy algorithm for goal-conditioned RL. Experiments demonstrate that, compared with prior RL methods, ours achieves 2×2$\times$2 × median improvement in success rates and can better cope with stochastic environments. In tabular settings, we show that our method is about 20×20$\times$20 × more sample efficient than the successor representation and 1500×1500$\times$1500 × more sample efficient than the standard (Monte Carlo) version of contrastive predictive coding.
@@ 1 Introduction
The main contribution of this paper is a temporal difference estimator for InfoNCE. We then apply this estimator to develop a new algorithm for goal-conditioned RL. Experiments on both state-based and image-based benchmarks show that our algorithm outperforms prior methods, especially on the most challenging tasks. Additional experiments demonstrate that our method can handle stochasticity in the environment more effectively than prior methods. We also demonstrate that our algorithm can be effectively applied in the offline setting. Additional tabular experiments demonstrate that TD InfoNCE is up to 1500×1500$\times$1500 × more sample efficient than the standard Monte Carlo version of the loss and that it can effectively stitch together pieces of data.
@@ 3.1 Preliminaries
[…]
@@ 3.2 Temporal Difference InfoNCE
$\mathcal{L}_{\text{TD InfoNCE}}(f)$ $\triangleq\mathbb{E}_{\begin{subarray}{c}(s,a)\sim p(s,a)\\ s_{t+}^{(2:N)}\sim p(s_{t+})\end{subarray}}\left[(1-\gamma)\mathbb{E}_{s_{t+}^{(1)}\sim p(s^{\prime}\mid s,a)}\left[\log\frac{e^{f(s,a,s_{t+}^{(1)})}}{\sum_{i=1}^{N}e^{f(s,a,s_{t+}^{(i)})}}\right]\right.$ (18)
@@ 4.1 Comparing to Prior Goal-conditioned RL methods
We report results in Fig. 1(a), and defer the full learning curves to Appendix Fig. 6. These results show that TD InfoNCE matches or outperforms other baselines on all tasks, both for state and image observations. On those more challenging tasks (pick & place (state / image) and slide (state / image)), TD InfoNCE achieves a 2×2$\times$2 × median improvement relative to the strongest baseline (Appendix Fig. 6). On the most challenging tasks, image-based pick & place and slide, TD InfoNCE is the only method achieving non-negligible success rates. For those tasks where the success rate fails to separate different methods significantly (e.g., slide (state) and push (image)), we include comparisons using minimum distances of the gripper or the object to the goal over an episode in Appendix Fig. 7. We speculate this observation is because TD InfoNCE estimates the discounted state occupancy measure more accurately, a hypothesis we will investigate in Sec. 4.3.
@@ 4.3 Accuracy of the estimated discounted state occupancy measure
Figure 3: Estimating the discounted state occupancy measure in a tabular setting. (Left) Temporal difference methods have lower errors than the Monte Carlo method. Also note that our TD InfoNCE converges as fast as the best baseline (successor representation). (Right) TD InfoNCE is more data efficient than other methods. Using a dataset of size 10M, TD InfoNCE achieves an error rate $25\%$ lower than the best baseline; TD InfoNCE also matches the performance of C-learning with 130×130$\times$130 × less data.
@@
Our next experiments studies sample efficiency. We hypothesize that the softmax in the TD InfoNCE loss may provide more learning signal than alternative methods, allowing it to achieve lower error on a fixed budget of data. To test this hypothesis, we run experiments with dataset sizes from 1K to 10M on the same gridworld, comparing TD InfoNCE to the same set of baselines. We report results in Fig. 3 with errors showing one standard deviation after training for 50K gradient steps for each approach. These results suggest that methods based on temporal difference learning predict more accurately than Monte Carlo method when provided with the same amount of data. Compared with its Monte Carlo counterpart, TD InfoNCE is 1500×1500$\times$1500 × more sample efficient ($6.5\times 10^{3}$ vs $10^{7}$ transitions). Compared with the only other TD method applicable in continuous settings (C-learning), TD InfoNCE can achieve a comparable loss with 130×130$\times$130 × less data ($7.7\times 10^{4}$ vs $10^{7}$ transitions). Even compared with the strongest baseline (successor representations), which makes assumptions (tabular MDPs) that our method avoids, TD InfoNCE can achieve a comparable error rate with almost 20×20$\times$20 × fewer samples ($5.2\times 10^{5}$ vs $10^{7}$ transitions).
@@ 4.4 Does TD InfoNCE enable off-policy reasoning?
Refer to caption Dataset Refer to caption TD InfoNCE Refer to caption Contrastive RL
Refer to caption Dataset Refer to caption TD InfoNCE Refer to caption Contrastive RL
@@ Appendix A Theoretical Analysis
¯$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)$ $\triangleq\mathbb{E}_{p(s,a)}\left[(1-\gamma)\mathbb{E}_{p(s^{\prime}=s_{t+}\mid s,a)}\left[\log C(s,a,s_{t+})\right]\right.$
@@
This objective is similar to ${\mathcal{L}}_{\text{TD InfoNCE}}$, but differs in that (a) softmax functions are replaced by $C(s,a,s_{t+})$ up to constant $\frac{1}{N\cdot p(s_{t+})}$ and (b) $w(s^{\prime},a^{\prime},s_{t+}^{(1:N)})$ is replaced by $\frac{C(s^{\prime},a^{\prime},s_{t+})}{p(s_{t+})}$. Formally, ${\mathcal{L}}_{\text{TD InfoNCE}}(C)$ is a nested Monte Carlo estimator of ¯$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}$ (Rainforth et al., 2018; Giles, 2015) and we leave the analysis of the gap between them as future works. We now find the solution of ¯$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)$ analytically by rewriting it using the cross entropy and ignore the stop gradient operator to reduce clutter: ¯$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)=$
$\mathbb{E}_{p(s,a)}\left[(1-\gamma)\mathbb{E}_{p(s^{\prime}=s_{t+}\mid s,a)}\left[\log C(s,a,s_{t+})\right]+\gamma\mathbb{E}_{\begin{subarray}{c}p(s^{\prime}\mid s,a),\pi(a^{\prime}\mid s^{\prime},g)\\ C(s^{\prime},a^{\prime},s_{t+})\end{subarray}}\left[\log C(s,a,s_{t+})\right]\right]$ (31)
@@
by definition. Since the (expected) softmax function is invariant to translation, we can write $f^{\star}(s,a,s_{t+})=\log p^{\pi}(s_{t+}\mid s,a)-\log p(s_{t+})-\log c(s,a)$, where $c(s,a)$ is an arbitrary function that does not depend on $s_{t+}$  [^2] Technically, $f^{\star}$ should be a set of functions satisfying $\left\{f:\frac{e^{f(s,a,s_{t+})}}{\mathbb{E}_{p(s_{t+}^{\prime})}\left[e^{f(s,a,s_{t+}^{\prime})}\right]}=\frac{C^{\star}(s,a,s_{t+})}{p(s_{t+})}\right\}.$. Thus, we conclude that TD InfoNCE objective converges to the same solution as that of MC InfoNCE (Eq. 13), i.e. ¯$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(f^{\star})={\mathcal{L}}_{\text{MC InfoNCE}}(f^{\star})$.
@@ D.3 Offline Goal-conditioned RL Experiments
As shown in Table 1, TD InfoNCE matches or outperforms all baselines on 5 / 6 tasks. On tasks (medium-play-v2 and medium-diverse-v2), TD InfoNCE achieves a $+13\%$ improvement over contrastive RL, showing the advantage of temporal difference learning over the Monte Carlo approach with a fixed dataset. We conjecture that this benefit comes from the dynamic programming property of the TD method and will investigate this property further in later experiments (Sec. 4.4). Additionally, TD InfoNCE performs 1.4×1.4$\times$1.4 × better than GCBC and retains a 3.8×3.8$\times$3.8 × higher scores than DT on average, where these baselines use (autoregressive) supervised losses instead of TD learning. These results suggest that TD InfoNCE is also a competitive goal-conditioned RL algorithm in the offline setting.
@@ D.4 Off-Policy Reasoning Experiments
Our second set of experiments aim to compare the performance of TD InfoNCE against contrastive RL on searching shortcuts in skewed datasets. To study this, we collect different datasets of size 20K with trajectories conditioned on the same pair of initial state and goal, with $95\%$ of the time along a long path and $5$% of the time along a short path. Using these skewed datasets, we again train both methods for 50K gradient steps and then evaluate the policy performance on reaching the same goal starting from the same state. We show the goal-conditioned policies learned by the two approaches in Fig. 9. The observation that TD InfoNCE learns to take shortcuts even though those data are rarely seen, while contrastive RL follows the long paths dominating the entire dataset, demonstrates the advantage of temporal difference learning over its Monte Carlo counterpart in improving data efficiency.
@@ D.5 Implementations and Hyperparameters
Hyperparameters
Values
actor learning rate
$5\times 10^{-5}$
critic learning rate
$3\times 10^{-4}$
using $\ell_{2}$ normalized representations
yes
hidden layers sizes (for both actor and representations)
$(512,512,512,512)$
contrastive representation dimensions
$16$
@@
Hyperparameters
Values
batch size (on large- tasks)
256 $\to$ 1024
hidden layers sizes (for both actor and representations on large- tasks)
(512, 512, 512, 512) $\to$ (1024, 1024, 1024, 1024)
behavioral cloning regularizer coefficient $\lambda$
$0.1$
goals for actor loss
random states $\to$ future states
@@ E.1 Hyperparameter Ablations
We conduct ablation experiments to study the effect of different hyperparameters in Table 2, aiming to find the best hyperparameters for TD InfoNCE. For each hyperparameter, we selected a set of different values and conducted experiments on push (state) and slide (state), one easy task and one challenging task, for five random seeds. We report mean and standard deviation of success rates in Fig. 10. These results suggest that while some values of the hyperparameter have similar effects, e.g. actor learning rate $=5\times 10^{-5}$ vs $1\times 10^{-4}$, our choice of combination is optimal for TD InfoNCE.
@@ E.4 Representation Interpolation
$\frac{\sin(1-\alpha)\eta}{\sin\eta}\phi(s_{0},a_{\text{no-op}},g)+\frac{\sin\alpha\eta}{\sin\eta}\phi(g,a_{\text{no-op}},a),$

### v3

@@ Abstract
Predicting and reasoning about the future lie at the heart of many time-series questions. For example, goal-conditioned reinforcement learning can be viewed as learning representations to predict which states are likely to be visited in the future. While prior methods have used contrastive predictive coding to model time series data, learning representations that encode long-term dependencies usually requires large amounts of data. In this paper, we introduce a temporal difference version of contrastive predictive coding that stitches together pieces of different time series data to decrease the amount of data required to learn predictions of future events. We apply this representation learning method to derive an off-policy algorithm for goal-conditioned RL. Experiments demonstrate that, compared with prior RL methods, ours achieves $2\times$ median improvement in success rates and can better cope with stochastic environments. In tabular settings, we show that our method is about $20\times$ more sample efficient than the successor representation and $1500\times$ more sample efficient than the standard (Monte Carlo) version of contrastive predictive coding.
@@ 1 Introduction
The main contribution of this paper is a temporal difference estimator for InfoNCE. We then apply this estimator to develop a new algorithm for goal-conditioned RL. Experiments on both state-based and image-based benchmarks show that our algorithm outperforms prior methods, especially on the most challenging tasks. Additional experiments demonstrate that our method can handle stochasticity in the environment more effectively than prior methods. We also demonstrate that our algorithm can be effectively applied in the offline setting. Additional tabular experiments demonstrate that TD InfoNCE is up to $1500\times$ more sample efficient than the standard Monte Carlo version of the loss and that it can effectively stitch together pieces of data.
@@ 3.1 Preliminaries
$\mathbb{E}_{p(y)}\left[\exp\left(f^{\star}(x,y)\right)\right]=\int\cancel{p(y)}\frac{p(y\mid x)}{\cancel{p(y)}c(x)}dy=\frac{1}{c(x)}\underbrace{\int p(y\mid x)dy}_{=1}=\frac{1}{c(x)}.$ (2)
@@ 3.2 Temporal Difference InfoNCE
$\mathcal{L}_{\text{TD InfoNCE}}(f)$ $\triangleq\mathbb{E}_{\begin{subarray}{c}(s,a)\sim p(s,a)\\ s_{t+}^{(2:N)}\sim p(s_{t+})\end{subarray}}\left[(1-\gamma)\mathbb{E}_{s_{t+}^{(1)}\sim p(s^{\prime}\mid s,a)}\left[\log\frac{e^{f(s,a,s_{t+}^{(1)})}}{\sum_{i=1}^{N}e^{f(s,a,s_{t+}^{(i)})}}\right]\right$.
@@ 4.1 Comparing to Prior Goal-conditioned RL methods
We report results in Fig. 2(a), and defer the full learning curves to Appendix Fig. 6. These results show that TD InfoNCE matches or outperforms other baselines on all tasks, both for state and image observations. On those more challenging tasks (pick & place (state / image) and slide (state / image)), TD InfoNCE achieves a $2\times$ median improvement relative to the strongest baseline (Appendix Fig. 6). On the most challenging tasks, image-based pick & place and slide, TD InfoNCE is the only method achieving non-negligible success rates. For those tasks where the success rate fails to separate different methods significantly (e.g., slide (state) and push (image)), we include comparisons using minimum distances of the gripper or the object to the goal over an episode in Appendix Fig. 7. We speculate this observation is because TD InfoNCE estimates the discounted state occupancy measure more accurately, a hypothesis we will investigate in Sec. 4.3.
@@ 4.3 Accuracy of the estimated discounted state occupancy measure
Figure 3: Estimating the discounted state occupancy measure in a tabular setting. (Left) Temporal difference methods have lower errors than the Monte Carlo method. Also note that our TD InfoNCE converges as fast as the best baseline (successor representation). (Right) TD InfoNCE is more data efficient than other methods. Using a dataset of size 10M, TD InfoNCE achieves an error rate $25\%$ lower than the best baseline; TD InfoNCE also matches the performance of C-learning with $130\times$ less data.
@@
Our next experiments studies sample efficiency. We hypothesize that the softmax in the TD InfoNCE loss may provide more learning signal than alternative methods, allowing it to achieve lower error on a fixed budget of data. To test this hypothesis, we run experiments with dataset sizes from 1K to 10M on the same gridworld, comparing TD InfoNCE to the same set of baselines. We report results in Fig. 3 with errors showing one standard deviation after training for 50K gradient steps for each approach. These results suggest that methods based on temporal difference learning predict more accurately than Monte Carlo method when provided with the same amount of data. Compared with its Monte Carlo counterpart, TD InfoNCE is $1500\times$ more sample efficient ($6.5\times 10^{3}$ vs $10^{7}$ transitions). Compared with the only other TD method applicable in continuous settings (C-learning), TD InfoNCE can achieve a comparable loss with $130\times$ less data ($7.7\times 10^{4}$ vs $10^{7}$ transitions). Even compared with the strongest baseline (successor representations), which makes assumptions (tabular MDPs) that our method avoids, TD InfoNCE can achieve a comparable error rate with almost $20\times$ fewer samples ($5.2\times 10^{5}$ vs $10^{7}$ transitions).
@@ 4.4 Does TD InfoNCE enable off-policy reasoning?
Dataset
TD InfoNCE
Contrastive RL
@@ Appendix A Theoretical Analysis
$\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)$ $\triangleq\mathbb{E}_{p(s,a)}\left[(1-\gamma)\mathbb{E}_{p(s^{\prime}=s_{t+}\mid s,a)}\left[\log C(s,a,s_{t+})\right]\right$.
@@
This objective is similar to ${\mathcal{L}}_{\text{TD InfoNCE}}$, but differs in that (a) softmax functions are replaced by $C(s,a,s_{t+})$ up to constant $\frac{1}{N\cdot p(s_{t+})}$ and (b) $w(s^{\prime},a^{\prime},s_{t+}^{(1:N)})$ is replaced by $\frac{C(s^{\prime},a^{\prime},s_{t+})}{p(s_{t+})}$. Formally, ${\mathcal{L}}_{\text{TD InfoNCE}}(C)$ is a nested Monte Carlo estimator of $\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}$ (Rainforth et al., 2018; Giles, 2015) and we leave the analysis of the gap between them as future works. We now find the solution of $\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)$ analytically by rewriting it using the cross entropy and ignore the stop gradient operator to reduce clutter: $\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(C)=$
$\mathbb{E}_{p(s,a)}\left[(1-\gamma)\mathbb{E}_{p(s^{\prime}=s_{t+}\mid s,a)}\left[\log C(s,a,s_{t+})\right]+\gamma\mathbb{E}_{\begin{subarray}{c}p(s^{\prime}\mid s,a),\pi(a^{\prime}\mid s^{\prime},g)\\ C(s^{\prime},a^{\prime},s_{t+})\end{subarray}}\left[\log C(s,a,s_{t+})\right]\right]$
@@
by definition. Since the (expected) softmax function is invariant to translation, we can write $f^{\star}(s,a,s_{t+})=\log p^{\pi}(s_{t+}\mid s,a)-\log p(s_{t+})-\log c(s,a)$, where $c(s,a)$ is an arbitrary function that does not depend on $s_{t+}$  [^2] Technically, $f^{\star}$ should be a set of functions satisfying $\left\{f:\frac{e^{f(s,a,s_{t+})}}{\mathbb{E}_{p(s_{t+}^{\prime})}\left[e^{f(s,a,s_{t+}^{\prime})}\right]}=\frac{C^{\star}(s,a,s_{t+})}{p(s_{t+})}\right\}.$. Thus, we conclude that TD InfoNCE objective converges to the same solution as that of MC InfoNCE (Eq. 7), i.e. $\bar{{\mathcal{L}}}_{\text{TD InfoNCE}}(f^{\star})={\mathcal{L}}_{\text{MC InfoNCE}}(f^{\star})$.
@@ D.3 Offline Goal-conditioned RL Experiments
As shown in Table 1, TD InfoNCE matches or outperforms all baselines on 5 / 6 tasks. On tasks (medium-play-v2 and medium-diverse-v2), TD InfoNCE achieves a $13\%$ improvement over contrastive RL, showing the advantage of temporal difference learning over the Monte Carlo approach with a fixed dataset. We conjecture that this benefit comes from the dynamic programming property of the TD method and will investigate this property further in later experiments (Sec. 4.4). Additionally, TD InfoNCE performs $1.4\times$ better than GCBC and retains a $3.8\times$ higher scores than DT on average, where these baselines use (autoregressive) supervised losses instead of TD learning. These results suggest that TD InfoNCE is also a competitive goal-conditioned RL algorithm in the offline setting.
@@ D.4 Off-Policy Reasoning Experiments
Our second set of experiments aim to compare the performance of TD InfoNCE against contrastive RL on searching shortcuts in skewed datasets. To study this, we collect different datasets of size 20K with trajectories conditioned on the same pair of initial state and goal, with $95\%$ of the time along a long path and 55% of the time along a short path. Using these skewed datasets, we again train both methods for 50K gradient steps and then evaluate the policy performance on reaching the same goal starting from the same state. We show the goal-conditioned policies learned by the two approaches in Fig. 9. The observation that TD InfoNCE learns to take shortcuts even though those data are rarely seen, while contrastive RL follows the long paths dominating the entire dataset, demonstrates the advantage of temporal difference learning over its Monte Carlo counterpart in improving data efficiency.
@@ D.5 Implementations and Hyperparameters
Hyperparameters Values
actor learning rate $5\times 10^{-5}$
critic learning rate $3\times 10^{-4}$
using $\ell_{2}$ normalized representations yes
hidden layers sizes (for both actor and representations) $(512,512,512,512)$
contrastive representation dimensions $16$
@@
Hyperparameters Values
batch size (on large- tasks) 256 $\to$ 1024
hidden layers sizes (for both actor and representations on large- tasks) (512, 512, 512, 512) $\to$ (1024, 1024, 1024, 1024)
behavioral cloning regularizer coefficient $\lambda$ $0.1$
goals for actor loss random states $\to$ future states
@@ E.1 Hyperparameter Ablations
We conduct ablation experiments to study the effect of different hyperparameters in Table 2, aiming to find the best hyperparameters for TD InfoNCE. For each hyperparameter, we selected a set of different values and conducted experiments on push (state) and slide (state), one easy task and one challenging task, for five random seeds. We report mean and standard deviation of success rates in Fig. 10. These results suggest that while some values of the hyperparameter have similar effects, e.g. actor learning rate $5\times 10^{-5}$ vs $1\times 10^{-4}$, our choice of combination is optimal for TD InfoNCE.
@@ E.4 Representation Interpolation
$\frac{\sin(1-\alpha)\eta}{\sin\eta}\phi(s_{0},a_{\text{no-op}},g)+\frac{\sin\alpha\eta}{\sin\eta}\phi(g,a_{\text{no-op}},g),$